OpenStudy (anonymous):

can someone please help me try and figure out where I'm going wrong with this definite integral? :)

OpenStudy (anonymous):

\[2\pi \int\limits_{0}^{4}x \sqrt(1+\frac{ x^2 }{ 16 }) dx\]

OpenStudy (anonymous):

Looks like an easy u sub.

OpenStudy (anonymous):

this was what I did: let u=1 + (X^2)16, therefore du= x/8 dx and dx=8/x du

OpenStudy (anonymous):

whoops typo divided by 16

OpenStudy (anonymous):

so then I got \[16\pi \int\limits_{0}^{4}\sqrt(u)du\]

OpenStudy (anonymous):

is that right so far?

OpenStudy (anonymous):

You should change the limits of integration.

OpenStudy (anonymous):

\[ (0,1)\to (u(0),u(1)) \]

OpenStudy (anonymous):

I can't carry it through to the end? That's the way we've done it in class. Shouldn't it give the same answer eventually?

OpenStudy (anonymous):

so for example I got \[\frac{ 2 }{ 3 }u ^{3/2} \] from 0 to 4, and then... finally I got

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }(1+\frac{ x^2 }{ 16 })\] evaluated from 0 to 4, and then that answer multiplied by \[16 \pi\]

OpenStudy (anonymous):

You're not getting it...

OpenStudy (anonymous):

When you do u sub, the integration limits change

OpenStudy (anonymous):

you can keep the same limits if you substitute whatever you substituted back in at the end, which I did though

OpenStudy (anonymous):

you cannot integrate with respect to u and use limits defined for x

OpenStudy (anonymous):

whatever expression you used to define your u, plug in x limits to get u limits

OpenStudy (anonymous):

yeah, but I plugged the \[1+\frac{ x^2 }{ 16 }\] back in at the end before I did my calculations, so I can't keep the limits?

OpenStudy (anonymous):

if x limits are 0 and 4, then for u(x)=1+x^2/16 u(0)=1+(0)^2/16=1 u(4)=1+(4)^2/16=2

OpenStudy (anonymous):

okay, just a sec let me see if I get the right answer using that :)

OpenStudy (anonymous):

your new limits of integration with respect to u are now 1 and 2, once you make the u substitution you can forget about x since you redifined your variable and therefore you need to redifine your limits

OpenStudy (anonymous):

I did. Thanks :)

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

make sure you understand this, because it will be on a test, 100% guaranteed